Herschel Status Report - June 2010
The spacecraft is operating nominally.
Star tracker warm pixels
Despite extensive preparations and tests that were performed successfully on ground with the avionics model (AVM), the upload to the spacecraft of a star tracker "bad pixel table" failed. After permanent reduction of the star tracker CCD temperature three months ago, this table, populated with the coordinates of pixels that should be ignored in the on-board attitude determination, is one more line of defence against a slowly increasing number of "warm" pixels. At present, failure to upload this table has no impact on the mission operations or science. Industry has worked out a procedure how the CCD temperature could be lowered further should this be deemed necessary. (See the previous status reports for the introduction of these measures to tackle the "warm" pixels).
With the exception of two single event upsets (SEU) in the HIFI memory, all three instruments continue to operate nominally. One of the HIFI SEUs manifested itself in a hitherto unknown way, overloading the spacecraft systems with corrupted telemetry. The only plausible scenario that has been identified to explain the anomaly is that on this occasion it was not the Local Oscillator Control Unit that was affected, but the Instrument Control Unit. The recovery activities were interrupted by another SEU, such that a total of almost three operational days for taking science data were lost.
The statistics of SEUs improve with the increase in elapsed mission time; it is now just over a year since Herschel's launch on 14 May 2009. In the satellite's solid state mass memory (SSMM), which is protected by error detection and correction (EDAC) circuitry, on the order of 1000 bit flips have been detected and reported since launch, none of which had any operational impact. In HIFI we have seen 11 bit flips in five months since operations were resumed using the redundant signal chain. Analysis of the SPIRE data show only 2 bit flips in one year. PACS has not had a single bit flip. All attempts made so far have failed to correlate the observed bit flips to any satellite environmental parameters, or to establish a temporal correlation with the occurrence of bit flips on the Planck spacecraft that also orbits about L2.
During the reporting period, mission operations have been conducted with the support of the ESA New Norcia ground station. Observational data stored on-board Herschel was received on ground during communication passes lasting approximately three hours.
The spacecraft's large Lissajous orbit about L2, which is perturbed by the helium outgassing and the daily reaction wheel biasing/unloading, has proven to be extremely stable. The last two planned orbit correction manoeuvres could be skipped because the required delta-v was below the threshold of 5 cm/s.
The ground segment is operating nominally. Data products are generated routinely and ingested into the Herschel Science Archive (HSA). Re-processing of all observations with version 2.1 of the data product pipeline has been completed successfully.
- 22 July 2010: Closing date of first in-flight Announcement of Opportunity (AO) for Open Time (OT1) observations
- 1 November 2010: Announcement of approved OT1 proposals